# FineMolTex: Towards Fine-grained Molecular Graph-Text Pre-training

### v1

@@ 3.1 Key Components of FineMolTex
Graph Encoder. We utilize GraphMVP [25], a pre-trained Graph Isomorphism Network (GIN), to obtain the embedding of each motif token. It employs multi-view pre-training to connect the 2D topologies and 3D geometries in the GEOM dataset [2] containing 250K conformations. Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ denote a motif with $N$ atoms, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is the atom set, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the bond set, and $\mathbf{X}=[\mathbf{x_{1},x_{2},\dots,x_{N}}]\in\mathbb{R}^{N\times\zeta}$ represents the atom feature matrix, where $\mathbf{x_{i}}$ is the feature vector of atom $v_{i}$ and $\zeta$ is the dimension of atom features. The embedding of each atom in the $l$-th layer is formed as
$\mathbf{g_{v}}^{(l)}=\text{COMBINE}^{(l)}\left(\mathbf{g_{v}}^{(l-1)},\text{AGGREGATE}^{(l)}\left(\{\mathbf{g_{u}}^{(l-1)}:u\in\mathcal{N}(v)\}\right)\right),$ (1)
where $\mathbf{g_{v}}^{(l)}$ is the embedding of atom $v$ at $l$-th layer, $\mathcal{N}(v)$ is the neighbors of atom $v$, AGGREGATE(·) aggregates the neighborhood information of atom $v$, and COMBINE(·) further combines the self-information with the neighborhood information. Then, we pool the atom embeddings at the final layer $\ell_{\text{g}}$ into a motif-level embedding, $\mathbf{g_{\mathcal{G}}}$, as follows.
$\mathbf{g_{\mathcal{G}}}=\text{READOUT}(\{\mathbf{g_{v}^{(\ell_{\text{g}})}}|v\in\mathcal{G}\}),\quad\text{for}\quad\mathcal{G}\in\{m_{1},m_{2},\dots,m_{J}\},$ (2)
@@
Text Encoder. We use SciBERT [3], which has been pre-trained on texts from the chemical and biological domains, as our text encoder, denoted as $f_{\text{bert}}$. It can encode a text sequence as:
@@
$\mathbf{z_{t_{0}},z_{t_{1}},\dots,z_{t_{D}}}$ $=f_{\text{trm}_{T}}(\mathbf{b_{t_{0}},b_{t_{1}},\dots,b_{t_{D}}}),$ (4)
$\mathbf{z_{m_{0}},z_{m_{1}},\dots,z_{m_{J}}}$ $=f_{\text{trm}_{M}}(\mathbf{g_{m_{0}},g_{m_{1}},\dots,g_{m_{J}}}).$ (5)
@@ 4.5 Ablation Study for Masking and Cross-Attention Layers (RQ5)
To thoroughly explore the impact of the key components in FineMolTex, we compare to several variants, including FineMolTex w/o mmm, which drops the masked multimodal modeling task altogether; FineMolTex w/o mask word, which does not mask word tokens; FineMolTex w/o mask motif, which does not mask motif tokens; FineMolTex w/o cross, which excludes cross-attention layers. We evaluate these variants on the graph-text retrieval task used in RQ1, on two datasets with $T=4$. As reported in Table LABEL:tab:abla, our model consistently surpasses the other variants. Specifically, without the masked multimodal modeling task, “FineMolTex w/o mmm” fails to capture fine-grained knowledge, resulting in the poorest performance. While “FineMolTex w/o mask word” and “FineMolTex w/o mask motif” perform better than “FineMolTex w/o mmm,” they are less effective than FineMolTex, which masks both words and motifs for mutual alignment. Lastly, without the cross-attention layers, “FineMolTex w/o cross” cannot integrate token embeddings from different modalities, thereby hampering its ability to effectively learn fine-grained knowledge. These observations demonstrate the effectiveness of each component.
@@
FineMolTex w/o mmm 70.33±$\pm 0.28$ 70.86±$\pm 0.16$ 92.56±$\pm 0.10$ 92.12±$\pm 0.25$
FineMolTex w/o mask word 72.85±$\pm 0.46$ 71.15±$\pm 0.24$ 93.45±$\pm 0.44$ 92.98±$\pm 0.18$
FineMolTex w/o mask motif 74.27±$\pm 0.07$ 72.11±$\pm 0.13$ 94.86±$\pm 0.32$ 93.88±$\pm 0.15$
FineMolTex w/o cross 71.44±$\pm 0.32$ 70.92±$\pm 0.18$ 92.66±$\pm 0.08$ 92.85 ±$\pm 0.24$

### v2

@@ 3.1 Key Components of FineMolTex
Graph Encoder. Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})$ represent a molecular graph with $N$ atoms, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of atoms, $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the bonds, and $\mathbf{X}=[\mathbf{x_{1},x_{2},\dots,x_{N}}]\in\mathbb{R}^{N\times\zeta}$ is the atom feature matrix. Here, $\mathbf{x_{i}}$ is the feature vector of atom $v_{i}$, and $\zeta$ is the dimension of atom features. We utilize GraphMVP [25], a pre-trained Graph Isomorphism Network (GIN), to encode each motif token. GraphMVP employs multi-view pre-training to connect 2D topologies and 3D geometries, leveraging the GEOM dataset [2], which contains 250K molecular conformations. Denoting the GraphMVP encoder as $f_{\text{GraphMVP}}$, we encode each atom $v$ into an embedding as follows:
$\mathbf{g_{v}}=f_{\text{GraphMVP}}(\mathbf{x_{v}},\mathbf{x_{u}}),\;u\in\mathcal{N}(v),$ (1)
where $\mathcal{N}(v)$ denotes the set of neighboring atoms of $v$. Then we pool the atom embeddings into a motif-level embedding, $\mathbf{g_{\mathcal{G}}}$, as follows:
$\mathbf{g_{\mathcal{G}}}=\text{READOUT}(\{\mathbf{g_{v}}|v\in\mathcal{G}\}),\quad\text{for}\quad\mathcal{G}\in\{m_{1},m_{2},\dots,m_{J}\},$ (2)
@@
To preserve the intrinsic connectivity of the motifs in the original molecule, we generate position embeddings based on the breadth-first search order and incorporate them into the motif embeddings $\mathbf{g_{m_{0}},g_{m_{1}},\dots,g_{m_{J}}}$, resulting in updated embeddings $\mathbf{g^{\prime}_{m_{0}},g^{\prime}_{m_{1}},\dots,g^{\prime}_{m_{J}}}$. Text Encoder. We use SciBERT [3], which has been pre-trained on texts from the chemical and biological domains, as our text encoder, denoted as $f_{\text{bert}}$. It can encode a text sequence as:
@@
Subsequently, we add the position embeddings to the token embeddings following previous work [3], yielding $\mathbf{b^{\prime}_{t_{0}},b^{\prime}_{t_{1}},\dots,b^{\prime}_{t_{D}}}$.
@@
$\mathbf{z_{t_{0}},z_{t_{1}},\dots,z_{t_{D}}}$ $=f_{\text{trm}_{T}}(\mathbf{b^{\prime}_{t_{0}},b^{\prime}_{t_{1}},\dots,b^{\prime}_{t_{D}}}),$ (4)
$\mathbf{z_{m_{0}},z_{m_{1}},\dots,z_{m_{J}}}$ $=f_{\text{trm}_{M}}(\mathbf{g^{\prime}_{m_{0}},g^{\prime}_{m_{1}},\dots,g^{\prime}_{m_{J}}}).$ (5)
@@ 4.5 Ablation Study for Masking and Cross-Attention Layers (RQ5)
To thoroughly explore the impact of the key components in FineMolTex, we compare to several variants, including w/o mmm, which drops the masked multimodal modeling task altogether; motif mask only, which only mask motif tokens; word mask only, which only mask word tokens; w/o cross-attention, which excludes cross-attention layers. We evaluate these variants on the graph-text retrieval task used in RQ1, on two datasets with $T=4$. As reported in Table LABEL:tab:abla, FineMolTex consistently surpasses the other variants. Specifically, without the masked multimodal modeling task, “w/o mmm” fails to capture fine-grained knowledge at all, resulting in the poorest performance. “motif mask only” and “word mask only” outperform “w/o mmm,” because they still enable some level of fine-grained knowledge learning by either predicting motifs based on word tokens or predicting word tokens based on motif tokens. However, they are less effective than FineMolTex, which masks both words and motifs for mutual alignment. Lastly, without the cross-attention layers, “w/o cross-attention” cannot integrate token embeddings from different modalities, thereby hampering its ability to effectively learn fine-grained knowledge. These observations demonstrate the effectiveness of each component.
@@
w/o mmm 70.33±$\pm 0.28$ 70.86±$\pm 0.16$ 92.56±$\pm 0.10$ 92.12±$\pm 0.25$
motif mask only 72.85±$\pm 0.46$ 71.15±$\pm 0.24$ 93.45±$\pm 0.44$ 92.98±$\pm 0.18$
word mask only 74.27±$\pm 0.07$ 72.11±$\pm 0.13$ 94.86±$\pm 0.32$ 93.88±$\pm 0.15$
w/o cross-attention 71.44±$\pm 0.32$ 70.92±$\pm 0.18$ 92.66±$\pm 0.08$ 92.85 ±$\pm 0.24$